OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

OpenStudy (nirmalnema):

area=side*side. =6.25*6.25 =39.0625 39.1

OpenStudy (anwaarullah):

6.25^2=39.1